10 squared plus 26 squared= C2- show me how to come up with the answer
Answer by tutorcecilia(2152)

10^2+26^2= C^2
100+676= C^2
776= C^2
sqrt(776)= sqrt(C^2) [Take the square root of both sides]
27.86=C
.
Check by plugging C=27.86 into the original equation:
10^2+26^2= C^2
10^2+26^2= 27.86^2
100 + 676=776.18
776=776.18 [Close enough]